### v1

# Gravity as a topological gauge theory

Abstract
We describe a topological gauge theory with underlying Cartan geometry $G/H$ and principal $H$-bundle. The topological action is constructed as a linear combination of the Euler and Pontrjagin numbers coming from entries of the curvature $\bar{\Omega}$ of a Cartan connection. As an example we show that for respectively $G=\left\{\begin{array}[]{ll}SO(4,1)\text{ for }\Lambda_{0}>0\\ SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ and $H=SO(3,1)$ a very particular linear combination yields the Holst + Euler and Pontrjagin of the curvature $R$ of the spin connection + Nieh-Yan + bare cosmological constant $\Lambda_{0}$ terms. Additionally, in this construction, the coupling constants of these different terms are inherently linked together. The other treated example is the one of the Möbius group $G=SO(4,2)/\{\pm I\}$ and $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$. By considering a constraint on the "pair of frames" $(\alpha,\beta)$ we retrieve an action comprising all terms compatible with diffeomorphism invariance and local Lorentz invariance described in the previous example + kinetic term for a scalar field $+$ an interaction of that scalar field with the Torsion $T$, the scalar field being related to dilation. Finally, in both examples, we study the equations of motion associated to the total (Gauge + Matter) action. It is especially shown that one of the solutions in the Möbius case consists of Einstein’s equations modified by the Holst and Nieh-Yan terms with an additional source term for curvature depending on specific variations of spin density of matter.
Keywords : Topological gravity, Cartan geometry, Characteristic classes, Conformal gravity
1 Introduction
There exist several approaches to describe gravity as a gauge theory. Two of them are MacDowell-Mansouri gravity and BF theory [17]. They can be mathematically described in terms of Cartan geometries $G/H$ of manifolds $\mathscr{M}$ where $G,H\subset G$ and $\mathfrak{g},\mathfrak{h}$ are respectively Lie groups and their associated Lie algebras. The associated $\mathfrak{g}$-valued Cartan connection $\varpi$ naturally splits into two parts $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$ when the geometry is reductive, one with values in $\mathfrak{h}$ and one in $\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. In these formalisms the $\mathfrak{h}$-valued part $\varpi_{\mathfrak{h}}$ corresponds to the spin connection $A$ while the $\mathfrak{p}$-valued part $\varpi_{\mathfrak{p}}=\sqrt{\dfrac{\Lambda}{3\epsilon}}\,\beta$ is proportional to the vielbein 1-form $\beta$ (tetrad if $dim(\mathscr{M})=4$). For a resulting Lorentzian [^1] Riemaniann or Galilean spacetimes can also be described by this formalism. For more details see [17, p10]. The kind of topological gauge theories developed in this article are also compatible with these choices of Lie groups. space-time the choice of Lie groups is given by $H=SO(3,1)$ while $G$ is chosen according to the value and sign of the cosmological constant $\Lambda$ as in [17, p10]:
$\left\{\begin{array}[]{lcll}G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda>0\\ G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda<0\end{array}\right.$ (3)
Definition 1.1.
Let $\mathscr{M}$ be a manifold, $G$ a Lie group and $H$ a closed subgroup of $G$ such that $dim(G/H)=dim(\mathscr{M})=m$ and $\mathscr{P}$ a principal $H$-bundle over $\mathscr{M}$. Let $\mathfrak{X}\in\mathfrak{h}$, we denote $\mathfrak{X}^{v}$ the associated vertical vector field on $\mathscr{P}$. A Cartan connection on $\mathscr{P}$ is equivalent to a $1$-form $\varpi$ on $\mathscr{P}$ with values in $\mathfrak{g}$ the Lie algebra of $G$ satisfying the conditions [9, p128]:
$(a)~{}\varpi(\mathfrak{X}^{v})=\mathfrak{X},~{}\text{ for any }\mathfrak{X}\in\mathfrak{h}.$ (4)
$(b)~{}(R_{h})^{*}\varpi=Ad(h^{-1})\varpi~{},\forall h\in H,\text{where }R\text{ is the right action of }H\text{ on }\mathscr{P}$ (5)
and $\quad\ \,\text{and }Ad\text{ the adjoint representation of }H\subset G\text{ on the Lie algebra }\mathfrak{g}.$
$(c)~{}\varpi(\mathfrak{X})\neq 0\text{ for every nonzero vector field }\mathfrak{X}\in\Gamma(T\mathscr{P}).$ (6)
Keywords : one, two, three, four
Let $\bar{\Omega}=d\varpi+\varpi\wedge\varpi=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{p}}$ be the curvature of the Cartan connection $\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}$ and $\star$ an internal Hodge as used in [12] such that $\star\kappa^{ab}=\tfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\kappa^{kl}$. For $m=4$, the MacDowell-Mansouri action is given by:
$S_{MM}[\varpi]$ $=-\dfrac{3}{2G\Lambda}\int_{\mathscr{M}}\operatorname{tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\star\bar{\Omega}_{\mathfrak{h}})=-\dfrac{3}{4G\Lambda}\int_{\mathscr{M}}\varepsilon_{abcd}\,\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}\ .$ (7)
This action is $Ad(H)$-invariant, i.e. Lorentz invariant. On the other hand, the Euler characteristic class $E(\mathscr{M})$ of a manifold $\mathscr{M}$ is null if $\dim(\mathscr{M})=m$ is odd, otherwise it can be linked to the $\tfrac{m}{2}$-th Pontrjagin class $P_{m/2}$ (more details can be found in [13]).
Definition 1.2.
In the case of an orientable differentiable manifold $\mathscr{M}$ of even dimension $m$, the Euler characteristic class can always  [^2] Since as indicated in [13] it is always possible to reduce the structure group of $T\mathscr{M}$ to $SO(m)$ via the Gram-Schmidt process. be expressed in terms of curvature as:
$E(\mathscr{M})=\operatorname{Pf}(\dfrac{\Omega_{\mathfrak{so}}}{2\pi})$
where $\operatorname{Pf}(\Omega_{\mathfrak{so}})$ is the Pfaffian of the $\mathfrak{so}(m)$-valued curvature 2-form $\Omega_{\mathfrak{so}}$. This is a $SO(m)$-invariant polynomial.
In the case of a $4$-dimensional manifold one has:
$E(\mathscr{M})=\tfrac{1}{2(4\pi)^{2}}\,\varepsilon_{abcd}\,\Omega_{\mathfrak{so}}^{ab}\wedge\Omega_{\mathfrak{so}}^{cd}$
More generally, let us give the
Definition 1.3.
The characteristic number $\mathdutchcal{C}$ associated to a given characteristic class $C(\mathscr{M})$ of $\mathscr{M}$ corresponds to the integral over $\mathscr{M}$ of its characteristic polynomial :
$\mathdutchcal{C}=\int_{\mathscr{M}}C(\mathscr{M}).$
In particular, for the case of the Euler class for $m=4$, the Euler number is thus:
$\mathdutchcal{E}=\dfrac{1}{2(4\pi)^{2}}\int_{\mathscr{M}}\varepsilon_{abcd}\,\Omega_{\mathfrak{so}}^{ab}\wedge\Omega_{\mathfrak{so}}^{cd}.$ (8)
Up to an overall factor, one can recognise the MacDowell-Mansouri action (7) is recast as the Euler number associated to the $\mathfrak{h}=\mathfrak{so}(3,1)$-valued part $\bar{\Omega}_{\mathfrak{h}}$ of the curvature and which is $SO(3,1$-invariant. This provides a very clear mathematical interpretation to the MacDowell-Mansouri action:
$S_{MM}[\varpi]=-\dfrac{24\pi^{2}}{G\Lambda}\,\mathdutchcal{E}$ (9)
Another characteristic class which will be of interest in the remainder of the article is the Pontrjagin characteristic class. As described in [13], one has the
Definition 1.4.
Let $L$ be a $k$-dimensional real vector bundle over an $m$-dimensional manifold $\mathscr{M}$. $\Omega$ is the curvature associated to a connection $\omega$ on the principal $H$-bundle over $\mathscr{M}$. Introducing orthonormal frames allows to reduce the structure group to $O(k)$ and one has $\mathfrak{o}(k)=\mathfrak{so}(k)$ at the level of Lie algebras. The $\mathfrak{so}(k)$-valued curvature is then $\Omega_{\mathfrak{so}}$. The corresponding total Pontrjagin class is then defined as: $P(\Omega_{\mathfrak{so}})=\det\Bigl{(}\mathds{1}+\dfrac{\Omega_{\mathfrak{so}}}{2\pi}\Big{)}$.
For a 4-dimensional manifold $\mathscr{M}$, the Pontrjagin number turns out to be:
$\mathdutchcal{P}=\int_{\mathscr{M}}P(\mathscr{M})=\int_{\mathscr{M}}P_{1}(\mathscr{M})=\int_{\mathscr{M}}-\frac{1}{8\pi^{2}}\operatorname{Tr}(\Omega_{\mathfrak{so}}\wedge\Omega_{\mathfrak{so}})\ ,$ (10)
where $P_{1}$ is the 1st Pontrjagin class. As stated in [13, Chern-Weil theorem], a nice feature of characteristic classes is that up to an exact term, they do not depend on the choice of connection.
Section 2 will introduce and describe the topological gauge theories associated to the Lie groups $G=SO(4,1),SO(3,2)$ considered as "mutations" of the groups $G=ISO(3,1)$. All these three groups contain $H=SO(3,1)$ which will be the structure group of the principal bundle $\mathscr{P}$. Then a topological action as a linear combination of the Euler and Pontrjagin numbers will be considered for these Lie groups $G$. Results of Section 2 are easily applicable to other spacetimes whether Riemaniann or Galilean or entirely different Lie groups. In particular, Section 3 will be devoted to the construction of a similar topological action for the case where $G=SO(4,2)/\{\pm I\}$ (the Möbius group). In the following, latin and greek letters will respectively be used for algebraic and space-time indices.
2 Characteristic polynomials for a Cartan connection with Lorentzian spacetimes
2.1 Lorentzian (dS and AdS) spacetimes as mutations of an
$ISO(3,1)/SO(3,1)$ Cartan geometry
Let $G$, $H\subset G$ be Lie groups with corresponding Lie algebras $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ such that $G/H$ is reductive, that is that the splitting $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{p}$ is $Ad(H)$-invariant. One such reductive model geometry is $G/H=ISO(3,1)/SO(3,1)$, with algebra $\mathfrak{g}=\overbrace{\mathfrak{so}(3,1)}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{\mathbb{R}}^{3,1}}^{\mathfrak{p}}$ where $G=ISO(3,1)=SO(3,1)\ltimes\mathbb{R}^{3,1}$ with $H=SO(3,1)$ defined by the relation $h^{\text{T}}\eta h=\eta~{},\forall h\in H$ where $\eta$ is the Minkowski metric:
$\eta=\begin{pmatrix}-1&0\\ 0&+I_{3}\end{pmatrix}\ .$
A group element $g\in G$ is given by :
$g=\begin{pmatrix}\Lambda&\dfrac{1}{\ell}v\\ 0&1\end{pmatrix}\in G\text{ where }\Lambda\in SO(3,1)\text{ such that }\Lambda^{\text{T}}\eta\Lambda=\eta\text{ and }v\in\mathbb{R}^{(3,1)}.$
Linearization leads to:
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\in\mathfrak{iso}(3,1)\text{ where }\varphi_{\mathfrak{h}}\in\mathfrak{so}(3,1)~{},~{}\varphi_{\mathfrak{h}}^{\text{T}}\eta+\eta\varphi_{\mathfrak{h}}=0~{},~{}\varphi_{\mathfrak{p}}=\dfrac{1}{\ell}\nu\in\mathbb{R}^{3,1}$
$=\begin{pmatrix}\frac{1}{2}\varphi^{ab}_{\mathfrak{h}}J_{ab}&\varphi^{a}_{\mathfrak{p}}e_{a}\\ 0&0\end{pmatrix}\in\mathfrak{iso}(3,1)$
with $\{J_{ab}\}_{a,b=0,...,3}$ generators of $\mathfrak{h}=\mathfrak{so}(3,1)$ and $\{e_{a}\}_{a=0,...,3}$ a basis of $\mathfrak{p}=\mathbb{R}^{3,1}$. In [16, p218], Sharpe introduces the notion of mutation of a Cartan geometry $G/H$ according to the
Definition 2.1.
Let $(\mathfrak{g},\mathfrak{h})$ and $(\mathfrak{g}^{\prime},\mathfrak{h})$ be two model geometries. A mutation map corresponds to an $Ad(H)$ module isomorphism $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ (i.e. $\mu(Ad_{h}(u))=Ad_{h}(\mu(u))\,$) satisfying:
$(i)~{}\mu_{|\mathfrak{h}}=\mathrm{id}_{\mathfrak{h}}$
$(ii)\bigl{[}\mu(u),\mu(v)\bigr{]}=\mu\bigl{(}[u,v]\bigr{)}\text{ mod }\mathfrak{h},\forall u,v\in\mathfrak{g}.$
The model geometry $(\mathfrak{g}^{\prime},\mathfrak{h})$ then corresponds to the mutant of the model geometry $(\mathfrak{g},\mathfrak{h})$ with the same group $H$.
A particular mutation that will be relevant for us is the one where $(\mathfrak{g},\mathfrak{h})=\bigr{(}\mathfrak{iso}(3,1),\mathfrak{so}(3,1)\bigr{)}$ and where the mutation map $\mu:\mathfrak{g}\rightarrow\mathfrak{g}^{\prime}$ is given by  [^3] More general mutations can be considered, e.g. $\varphi_{\mathfrak{g}}=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\longmapsto\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}\varphi_{\mathfrak{h}}&k\varphi_{\mathfrak{p}}\\ k^{\prime}\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}\in\mathfrak{g}^{\prime}$. However, whenever $k$ and $k^{\prime}$ are constants, this other mutation doesn’t change the physics that we will describe later on.
$\varphi_{\mathfrak{g}}=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ 0&0\end{pmatrix}\longrightarrow\varphi_{\mathfrak{g}^{\prime}}=\mu(\varphi_{\mathfrak{g}})=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}\in\mathfrak{g}^{\prime}$ (5)
with $k$ a 0-form scalar field and $\varphi_{\bar{\mathfrak{p}}}=\varphi_{\mathfrak{p}}^{\text{T}}\eta$. One can observe that $(\mathfrak{g}^{\prime},\mathfrak{h})$ reduces to dS and AdS geometries respectively for the two particular cases, see e.g. [17]:
$\left\{\begin{array}[]{lcll}k=-1\Leftrightarrow G=SO(4,1)&\text{ and }&\epsilon=+1,&\text{ for }\Lambda_{0}>0,\\ &&&\\ k=+1\Leftrightarrow G=SO(3,2)&\text{ and }&\epsilon=-1,&\text{ for }\Lambda_{0}<0.\end{array}\right.$ (9)
Throughout the paper $k$ will be considered as a real number, the case of a 00-form scalar field $k$ will be studied in a forthcoming paper. On the other hand, the canonical decomposition of a symmetric algebra $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$ (see [11, Prop.2.1, Chapter XI], and [10, chapter 11] for more examples) corresponds to the commutators:
$[\mathfrak{h},\mathfrak{h}]\subset\mathfrak{h},\qquad[\mathfrak{h},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{h}.$
Therefore, one can see that the mutated Lie algebra $\mathfrak{g}^{\prime}=\mathfrak{h}\oplus\mathfrak{m}$ of (5) corresponds to a symmetric algebra for $k\neq 0$, as a particular case of a reductive geometry with $\mathfrak{m}=\mathfrak{g}^{\prime}/\mathfrak{h}\simeq\mathfrak{p}=\mathfrak{g}/\mathfrak{h}$. Moreover, the generators of $\mathfrak{h}=\mathfrak{so}(3,1)$, and elements of the bases of $\mathfrak{p}=\mathfrak{\mathbb{R}}^{3,1}$, $\mathfrak{p}^{*}=\mathfrak{\mathbb{R}}^{*3,1}$ and ${\bar{\mathfrak{p}}}=\bar{\mathbb{R}}^{3,1}$ obey the relations (the isomorphism $\mathrm{End}(\mathbb{R}^{n})\simeq\mathbb{R}^{n}\otimes{\mathbb{R}^{n}}^{*}$ is used):
[…]
and
$J_{ab}e_{c}$ $=(J_{ab})^{r}_{~{}c}e_{r}=(\eta_{bc}\delta^{r}_{~{}a}-\eta_{ac}\delta^{r}_{~{}b})e_{r}=\eta_{bc}e_{a}-\eta_{ac}e_{b}$
$\bar{e}_{c}J_{ab}$ $=(\eta_{ac}\delta^{r}_{b}-\eta_{bc}\delta^{r}_{a})\bar{e}_{r}=\eta_{ac}\bar{e}_{b}-\eta_{bc}\bar{e}_{a}$
$[J_{ab},J_{cd}]$ $=\eta_{bc}J_{ad}+\eta_{ad}J_{bc}+\eta_{bd}J_{ca}+\eta_{ac}J_{db}=\tfrac{1}{2}C^{ef}_{ab,cd}J_{ef}$
with $C^{fe}_{ab,cd}=-C^{ef}_{ab,cd}$ which is easily computed. An element $\varphi=\varphi_{\mathfrak{h}}\oplus\varphi_{\mathfrak{m}}\in\mathfrak{g}^{\prime}$ admits the following matrix representation:
$\varphi$ $=\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}J_{ab}&\varphi_{\mathfrak{p}}^{a}e_{a}\\ k\varphi_{p}^{a}\bar{e}_{a}&0\end{pmatrix}=\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}\begin{pmatrix}J_{ab}&0\\ 0&0\end{pmatrix}\oplus\varphi_{\mathfrak{m}}^{a}\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ (18)
$=\dfrac{1}{2}\varphi_{\mathfrak{h}}^{ab}J_{ab}\oplus\varphi_{\mathfrak{m}}^{a}M_{a}$ (19)
with $\varphi^{a}_{\mathfrak{m}}=\varphi^{a}_{\mathfrak{p}}$ and $M_{a}=\begin{pmatrix}0&e_{a}\\ k\bar{e}_{a}&0\end{pmatrix}$ the generators of $\mathfrak{m}$. Let $\varphi,\varphi^{\prime}\in\mathfrak{g}^{\prime}$, their Lie bracket reads:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}\varphi_{\mathfrak{h}}&\varphi_{\mathfrak{p}}\\ k\varphi_{\bar{\mathfrak{p}}}&0\end{pmatrix},\begin{pmatrix}\varphi^{\prime}_{\mathfrak{h}}&\varphi^{\prime}_{\mathfrak{p}}\\ k\varphi^{\prime}_{\bar{\mathfrak{p}}}&0\end{pmatrix}\end{bmatrix}=\begin{pmatrix}[\varphi_{\mathfrak{h}},\varphi^{\prime}_{h}]+k(\varphi_{\mathfrak{p}}\varphi^{\prime}_{\bar{\mathfrak{p}}}-\varphi^{\prime}_{\mathfrak{p}}\varphi_{\bar{\mathfrak{p}}})&(\varphi_{\mathfrak{h}}\varphi^{\prime}_{\mathfrak{p}}-\varphi^{\prime}_{\mathfrak{h}}\varphi_{\mathfrak{p}})\\ k(\varphi_{\bar{\mathfrak{p}}}\varphi^{\prime}_{\mathfrak{h}}-\varphi^{\prime}_{\bar{\mathfrak{p}}}\varphi_{\mathfrak{h}})&0\end{pmatrix}$
$=\bigl{(}\dfrac{1}{8}\varphi_{\mathfrak{h}}^{ab}\varphi_{h}^{\prime cd}C_{ab,cd}^{ef}+\dfrac{k}{2}(\varphi_{\mathfrak{p}}^{e}\varphi_{\mathfrak{p}}^{\prime f}-\varphi_{\mathfrak{p}}^{f}\varphi_{\mathfrak{p}}^{\prime e})\bigr{)}J_{ef}\oplus\dfrac{1}{2}(\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{p}}^{\prime c}-\varphi_{\mathfrak{h}}^{\prime ab}\varphi_{\mathfrak{p}}^{c})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})M_{r}$ (20)
A Cartan connection $\varpi$ with this $G/H$ Cartan geometry and this mutation can be decomposed as:
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{m}}=\begin{pmatrix}\dfrac{1}{2}\varpi_{\mathfrak{h}}^{ab}J_{ab}&\varpi_{\mathfrak{p}}^{a}e_{a}\\ k\varpi_{\mathfrak{p}}^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}\dfrac{1}{2}A^{ab}J_{ab}&\dfrac{1}{\ell}\beta^{a}e_{a}\\ \dfrac{k}{\ell}\beta^{a}\bar{e}_{a}&0\end{pmatrix}=\begin{pmatrix}A&\dfrac{1}{\ell}\beta\\ \dfrac{k}{\ell}\bar{\beta}&0\end{pmatrix}$ (27)
where $\varpi_{\mathfrak{h}}^{ab}=\varpi^{ab}_{\mathfrak{h}\>\,\mu}dx^{\mu}=A^{ab}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the spin connection and $\ell\varpi_{\mathfrak{p}}^{a}=\beta^{\,a}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the tetrad while $\varpi_{\mathfrak{p}}=\dfrac{1}{\ell}\beta$ is the soldering form and $\bar{\beta}=\beta^{\text{T}}\eta$. The curvature associated to the covariant derivative and the Cartan connection is:
$\bar{\Omega}$ $=d\varpi+\dfrac{1}{2}[\varpi,\varpi]=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{m}}$ (28)
$=\bigl{(}d\varpi_{\mathfrak{h}}+\dfrac{1}{2}[\varpi_{\mathfrak{h}},\varpi_{\mathfrak{h}}]+k\xi\bigr{)}$ (29)
$\qquad\oplus\bigl{(}d\varpi_{\mathfrak{p}}+\dfrac{1}{4}(\varpi^{ab}_{\mathfrak{h}\>\,\mu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\nu}-\varpi^{ab}_{\mathfrak{h}\>\,\nu}\varpi^{~{}\,c}_{\mathfrak{p}\,\,\mu})(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})dx^{\mu}\wedge dx^{\nu}\bigr{)}\otimes M_{r}$
$=\bigl{(}\dfrac{1}{2}\partial_{\mu}A^{rs}_{~{}\,\,\nu}+\dfrac{1}{32}A^{ab}_{~{}\,\,\mu}A^{cd}_{~{}\,\,\nu}C_{ab,cd}^{rs}+\dfrac{k}{4}\xi^{rs}_{~{}~{}\mu\nu}\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes J_{rs}$
$\qquad\oplus\dfrac{1}{\ell}\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{4}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (30)
We identify the curvature and torsion of Einstein-Cartan gravity:
$R$ $=\bar{\Omega}_{\mathfrak{h}}-k\xi=dA+\dfrac{1}{2}[A,A]$ (31)
$T$ $=\ell\bar{\Omega}_{\mathfrak{m}}=d\beta+A\wedge\beta=\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{4}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (32)
with $\xi=\dfrac{1}{2}\xi^{ab}J_{ab}=\dfrac{1}{4}\xi^{ab}_{~{}~{}\mu\nu}dx^{\mu}\wedge dx^{\nu}\otimes J_{ab}$, where $\xi^{ab}=2\varpi_{\mathfrak{p}}^{a}\wedge\varpi_{\mathfrak{p}}^{b}$ and $\xi^{ab}_{~{}~{}\mu\nu}=2(\varpi^{~{}\,a}_{\mathfrak{p}\,\,\mu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\nu}-\varpi^{~{}\,a}_{\mathfrak{p}\,\,\nu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\mu})$. The Bianchi’s identities $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ split as:
[…]
$(D\bar{\Omega})_{\mathfrak{m}}$ $=\dfrac{1}{\ell}(dT+A\wedge T-R\wedge\beta)=0$
with $\bar{T}=T^{\text{T}}\eta$.
2.2 Hodge operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, the Killing form $K$ on $\mathfrak{g}$ splits into two parts:
[…]
In this case we choose $g=\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}}$. It corresponds to the pullback of the $\mathfrak{m}$ part of the Killing form on $\mathfrak{g}$ by $\varpi_{\mathfrak{m}}$ the $\mathfrak{m}$-part of the Cartan connection. For a manifold $\mathscr{M}$ of dimension $m=4$ with Lorentzian Cartan geometry defined by mutation (5):
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ $=K_{\mathfrak{h}}(\varphi_{h},\varphi^{\prime}_{h})+K_{\mathfrak{m}}(\varphi_{m},\varphi^{\prime}_{m})=3\operatorname{Tr}(\varphi_{h},\varphi^{\prime}_{h})+6k\eta(\varphi_{m},\varphi^{\prime}_{m})$
$=-3\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{h}^{\prime cd}+6k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{\prime b}$
with $K_{\mathfrak{m},ab}=6k\eta_{ab}$ and thus
$g(X,Y)=(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{6k}{\ell^{2}}\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}=\dfrac{6k}{\ell^{2}}\tilde{g}(X,Y)$
where $\tilde{g}=\beta^{*}\eta$ corresponds to the metric usually defined in the tetrad formalism. Let $\omega\in\Omega^{q}(\mathscr{M},\mathfrak{g})$ be a q-form on $\mathscr{M}$ with values in $\mathfrak{g}$. Its local trivialisation in a given chart is :
$\omega=\dfrac{1}{r!}\omega_{\mu_{1}\mu_{2}...\mu_{r}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{r}}$ (2)
Let $\varepsilon_{a_{1}a_{2}...a_{n}}$ be the Levi-Civita symbol of dimension n, such that $\varepsilon_{12...n}=1$. From the metric $g$ one can define a Hodge star operator *** that acts on $\omega$ as :
*$*\omega=$ $\dfrac{1}{r!}\sqrt{|det(g)|}\,\omega_{\mu_{1}...\mu_{r}}g^{\mu_{1}\nu_{1}}...g^{\mu_{r}\nu_{r}}\varepsilon_{\nu_{1}...\nu_{m}}dx^{\nu_{r+1}}\wedge...dx^{\nu_{m}}$ (3)
In order to stick with the standard literature, the volume form on $\mathscr{M}$ is denoted
$dvol=\sqrt{|\tilde{g}|}d^{m}x=\sqrt{|\tilde{g}|}\dfrac{1}{m!}\varepsilon_{\mu_{1}\mu_{2}...\mu_{m}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{m}}=\dfrac{1}{m!}\varepsilon_{a_{1}a_{2}...a_{m}}\beta^{\,a_{1}}\wedge\beta^{\,a_{2}}\wedge...\wedge\beta^{\,a_{m}}.$
2.3 Associated action
In the remainder of this article we relate the Levi-Civita symbol to the Levi-Civita tensor according to the relations :
$\varepsilon_{abcd}=\varepsilon_{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;b}\beta^{-1\,\mu_{3}}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\mu_{4}}_{~{}~{}\,\,\;\;\;d}$ and $\varepsilon^{abcd}=\varepsilon^{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{\,a}_{~{}\,\,\mu_{1}}\beta^{\,b}_{~{}\,\,\mu_{2}}\beta^{\,c}_{~{}\,\,\mu_{3}}\beta^{\,d}_{~{}\,\,\mu_{4}}$. The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the curvature $\bar{\Omega}$ having values in $\mathfrak{g}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+2k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}$. Taking a linear combination of the Euler and Pontrjagin numbers ($\mathdutchcal{E}$ (8) and $\mathdutchcal{P}$ (10)) associated respectively to $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}$ we can build the action :
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}ePf(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})+rP(\bar{\Omega})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}-\dfrac{r}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}\wedge\bar{\Omega})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+2kR^{ab}\wedge\xi^{cd}+k^{2}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+2kR^{ab}\wedge\xi_{ab}-\dfrac{2k}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{8k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{96k^{2}}{\ell^{4}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;d}\varepsilon^{cd}_{~{}~{}ab}d^{4}x-\dfrac{2k}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}.$ (1)
[…]
$S_{M}[A,\beta]$ $=\dfrac{\ell^{3}}{216k^{2}}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=ReS_{D}$ (2)
where $ReS_{D}$ corresponds to the real part of the Dirac action as described in [8, p.197]. We are now in position to identify the different terms of $S_{G}[A,\beta]$ above with the action presented in [14] which in our notations correspond to:
$S^{\prime}[\varpi]$ $=\int_{\mathscr{M}}\Biggl{(}\biggl{(}\overbrace{\dfrac{\alpha_{1}\ell^{2}}{2}R^{ab}\wedge\star\xi_{ab}+\dfrac{\alpha_{2}\ell^{2}}{2}R^{ab}\wedge\xi_{ab}}^{Holst}+\overbrace{\alpha_{3}R^{ab}\wedge R_{ab}}^{Pontrjagin}$ (3)
$+\overbrace{\alpha_{4}R^{ab}\wedge\star R_{ab}}^{Euler}+\overbrace{\alpha_{5}\bigl{(}T^{a}\wedge T_{a}-\dfrac{\ell^{2}}{2}R^{ab}\wedge\xi_{ab}\bigr{)}}^{Nieh-Yan}+\overbrace{\dfrac{\alpha_{6}\ell^{4}}{4}\varepsilon_{abcd}\xi^{ab}\wedge\xi^{cd}}^{Cosmological~{}constant}\biggr{)}\Biggr{)}$
where the $\star$ is an internal Hodge such that $\star\beta^{ab}=\dfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\beta^{kl}$ as in [12, p.2]. For $G=ISO(3,1)$ ($k=0$) the action simplifies to the topological terms:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}}^{Euler}+\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\Bigr{)}$ (4)
In this case from the first Bianchi identity $D\bar{\Omega}_{\mathfrak{h}}=0$, and by neglecting boundary terms, the equations of motion for $A$ and $\beta$ require respectively a null spin density and a null energy-momentum tensor and thus turns out to be physically irrelevant. While for general mutation $\mu$ with parameter $k$, (with particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (3) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell^{2}}=\alpha_{1}\\ \dfrac{rk}{2\pi^{2}\ell^{2}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell^{2}}=\alpha_{5}\\ \dfrac{ek^{2}}{8\pi^{2}\ell^{4}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell^{2}}\alpha_{3}=-2\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{r\ell^{2}}\alpha_{3}\\ r=\dfrac{2\pi^{2}\ell^{2}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (5)
The solution of (5) in terms of the parameters $e$, $r$, $k$, and $\ell$ is:
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{2k}{\ell^{2}}\alpha_{3}=-\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r}\alpha_{2}\\ r=\dfrac{4\pi^{2}\ell^{2}}{k}\alpha_{2}\end{cases}$ (6)
and the resulting action reads:
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}$ (7)
$=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell^{2}}\big{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+rR^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi)^{2}\ell^{4}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\quad+\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$ (8)
The equations of motion yield (neglecting boundary terms):
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}(R^{ab}+\dfrac{2k}{\ell^{2}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{\pi^{2}\ell^{2}}R^{ab}\wedge\beta_{b}\delta_{ac}=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta\beta^{c}}$ (9)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{\pi^{2}\ell^{2}}T_{a}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta A^{ab}}$ (10)
where $\tau$ and $\mathfrak{s}$ are respectively the energy-momentum tensor and the spin density of matter. For $r=0$ (no Pontrjagin term), setting $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$ and applying the Hodge star *** to the first equation (9) gives us the Einstein equations with bare cosmological constant $\Lambda_{0}$ while the other (10) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:
$G_{kc}+\Lambda_{0}\eta_{kc}$ $=-\dfrac{1}{2\alpha_{1}}\tau_{ck}$ (11)
$T^{c}\wedge\beta^{d}\varepsilon_{abcd}$ $=\dfrac{1}{2\alpha_{1}}\mathfrak{s}_{ab}\ .$ (12)
Separating the energy-momentum tensor $\tau_{ck}$ as $\tau_{ck}=\tau_{M,ck}-\rho_{\text{vac}}\eta_{ck}$ with $\tau_{\text{vac},ck}=-\rho_{\text{vac}}\eta_{ck}$ [4] the part related to the vacuum energy density $\rho_{\text{vac}}$ leads to:
$G_{kc}+\Lambda\eta_{kc}$ $=-\dfrac{1}{2\alpha_{1}}\tau_{M,ck}$ (13)
$T^{c}\wedge\beta^{d}\varepsilon_{abcd}$ $=\dfrac{1}{2\alpha_{1}}\mathfrak{s}_{ab}$ (14)
where $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell^{2}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell^{2}}-\dfrac{2\pi^{2}\ell^{2}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ is the effective cosmological constant, and $\Lambda_{\text{vac}}$ is the vacuum contribution. Let $\Lambda_{exp}$ be the measured value of the effective cosmological constant $\Lambda$ and several situations occur according to the theoretical predictions for $\rho_{\text{vac}}$.
  • If $\Lambda_{\text{vac}}>\Lambda_{exp}$, then, agreement with experiment could be attained by fixing the appropriate value of $\Lambda_{0}=-\dfrac{6k}{\ell^{2}}$ with $k>0$, in particular, for $k=1$ $\Leftrightarrow G=SO(3,2)$ (AdS spacetime), thus potentially allowing the use of AdS/CFT correspondence [1].
  • If $\Lambda_{\text{vac}}<\Lambda_{exp}$, agreement with $\Lambda_{exp}$ would require a mutation parameter $k<0$. A particular case is $k=-1$ $\Leftrightarrow G=SO(4,1)$ (dS spacetime). Especially, $\rho_{\text{vac}}=0$ whenever we impose normal ordering as mentionned in [5].
  • The last case is $\Lambda_{\text{vac}}=\Lambda_{exp}$ this implies $k=0\Leftrightarrow G=ISO(3,1)$, which corresponds to Minkoswki spacetime. However, for $k=0$ we do not retrieve Einstein’s equations since the action consists solely of the total derivatives corresponding to both the Euler and Pontryagin densities of the curvature $R$.
Remark.
The equations of motion are the same as those of Einstein-Cartan gravity modulo the boundary terms if $r=0$. However, even if they do not affect the classical equations of motion, the Euler, Pontrjagin and Nieh-Yan densities (which are not present in usual Einstein-Cartan gravity) could have non trivial effects in a path integral formulation of quantum gravity when summing over all possible configurations as stated in [14, p2].
For $r\neq 0$ we can identify $\alpha_{5}=-\alpha_{2}=-\dfrac{rk}{4\pi^{2}\ell^{2}}=-\dfrac{2\alpha_{1}}{3\gamma}$, with $\gamma$ the Barbero-Immirzi parameter. By replacing $e$, $r$ and $\ell$ in the action we obtain:
$S_{G}[\varpi]$ $=\dfrac{3\pi}{2\Lambda_{0}G}(\mathdutchcal{E}+\dfrac{2}{3\gamma}\mathdutchcal{P})$ (15)
$=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}$
Thus, by building an action from a linear combination of the Euler and Pontrjagin numbers associated to a manifold with Lorentzian Cartan geometry we retrieve the Palatini action with a bare cosmological constant $\Lambda_{0}$ as well as the Euler, Pontrjagin and Nieh-Yan densities of the curvature $R$ and torsion $T$. According to [6], these three last terms should be present in any quantum theory of gravity. The main difference when comparing with the litterature is that in our case the coupling constants of the Euler and Pontrjagin densities are determined by the cosmological constant and by respectively the coupling constant of the Palatini action and the Barbero-Immirzi parameter. These dependencies on $G,\gamma$, and $\Lambda_{0}$ are interesting since a quantum version of the theory may yield effects sensitive to the bare cosmological constant $\Lambda_{0}$ in particular. We believe that probing the bare cosmological constant this way could be highly beneficial in constraining the vacuum contribution to the cosmological constant $\Lambda_{\text{vac}}$ since the two cases $\Lambda_{\text{vac}}=0$ and $\Lambda_{\text{vac}}=8\pi G\rho_{vac}$ bring about huge differences in the numerical value of $\Lambda_{0}$. One can also see that the coupling constant of the Euler density $\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}$ obtained via the equation of motion (9) for the tetrad $\beta$ matches perfectly with what has been described in [2, section 3]. And in case $r=0$, it makes the action invariant under the duality symmetry consisting of swapping $R^{ab}\leftrightarrow\tfrac{\Lambda_{0}}{3}\beta^{a}\wedge\beta^{b}$. The case of a dynamical $\ell$ will be studied in a forthcoming paper. Also, in our case, properties of Characteristic classes imply that the choice of connection has no impact on the gauge action up to exact terms. We can consider $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ as fixed. Therefore, from the way we built the action, $r$ is the only remaining free parameter and is linked to the Barbero-Immirzi parameter $\gamma$ via the relation $r=\dfrac{\pi}{\Lambda_{0}G\gamma}$.
Remark.
The Pontrjagin density and other terms described here correspond to the Chern-Simons modified gravity [3] terms with a constant scalar field $\vartheta=\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}$ and potential term for $\vartheta$, $S_{\vartheta}=\int_{\mathscr{M}}V(\vartheta)$, with :
$V(\vartheta)$ $=\dfrac{2\alpha_{1}}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\varepsilon_{abcd}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}}^{Cosmological~{}constant}$ (16)
$\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{i}\wedge\beta_{j}\bigr{)}}^{Nieh-Yan}.$
The Chern-Simons term also occurs in Loop Quantum Gravity (LGQ) and String theory [3].
Let us collect some of the results in the table below:
$\left\{\begin{array}[]{ll}k=-1&\Leftrightarrow G=SO(4,1)\text{ for }\Lambda_{0}>0\\[5.69054pt] \left\{\begin{array}[]{ll}k=\pm 1\Leftrightarrow G=SO(4,1),SO(3,2)\\ \ell\rightarrow\infty\\ r=0\end{array}\right\}\Leftrightarrow\left\{\begin{array}[]{ll}\Lambda_{0}\rightarrow 0\\ \gamma\rightarrow\infty\end{array}\right\}&\simeq\text{Einstein-Cartan}\\ k=1&\Leftrightarrow G=SO(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$ (25)
Taking $r\neq 0$ in the action (7) (with Pontrjagin number $\mathdutchcal{P}$ relative to the Cartan geometry $G/H$) secures the presence of the Barbero-Immirzi part of the action, together with the Pontrjagin density of $R$ and the Nieh-Yan term.
3 Characteristic polynomials and action for the Möbius group $G=SO(4,2)/\{\pm I\}$
Another interesting example of geometry is given by the Möbius group $G=SO(4,2)/\{\pm I\}$ defined by the relation $g^{T}Ng=N$, $g\in G$, with
$N=\begin{pmatrix}0&0&-1\\ 0&\eta&0\\ -1&0&0\end{pmatrix}$
and $H=CO(4,2)\ltimes\mathbb{R}^{*3,1}$ the maximal normal subgroup of $G$. A group element of H admits the matrix representation :
$H\ni h=\overbrace{\begin{pmatrix}z&0&0\\ 0&\Lambda&0\\ 0&0&z^{-1}\end{pmatrix}}^{\in G_{0}=CO(4,2)}\overbrace{\begin{pmatrix}1&\bar{u}&\tfrac{1}{2}\bar{u}u\\ 0&\mathds{1}&u\\ 0&0&1\end{pmatrix}}^{\in G_{1}=\mathbb{R}^{*3,1}}\text{ where }\Lambda\in SO(3,1),\Lambda^{\text{T}}\eta\Lambda=\eta\;;\;z\Lambda\in CO(4,2)$
$u\in\mathbb{R}^{3,1}\;;\;\bar{u}=u^{\text{T}}\eta\in\mathbb{R}^{*3,1}\text{ and }z\in\mathbb{R}.$
Its Lie algebra is $\mathfrak{g}=\mathfrak{so}(4,2)=\mathfrak{g}_{-1}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{1}$ with $\mathfrak{g}_{-1}\simeq\mathbb{R}^{3,1}$, $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)=\mathfrak{so}(3,1)\oplus\mathbb{R}$ and $\mathfrak{g}_{1}\simeq\mathbb{R}^{*3,1}$, turns out to be a graded Lie algebra [9], namely,
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{g}_{-1}]\subset\mathfrak{g}_{-1},\qquad[\mathfrak{g}_{1},\mathfrak{g}_{0}]\subset\mathfrak{g}_{1},\qquad[\mathfrak{g}_{-1},\mathfrak{g}_{1}]\subset\mathfrak{g}_{0}$
$\qquad\qquad\qquad\qquad\qquad\;\;[\mathfrak{g}_{-1},\mathfrak{g}_{-1}]=[\mathfrak{g}_{1},\mathfrak{g}_{1}]=0.$
The generators of the above graded algebra are the same as the ones described previously in (2.1) except for $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ which is generated by the $\{J_{ab}\}$’s and $\mathds{1}_{4}$. An element of $\mathfrak{g}_{0}\simeq\mathfrak{co}(4,2)$ is $\gamma_{\mathfrak{co}}=\frac{1}{2}\gamma_{\mathfrak{so}}^{ab}J_{ab}-z\mathds{1}_{4}$ with $\gamma_{\mathfrak{so}}\in\mathfrak{so}(3,,1)$.
3.1 The Möbius group as a symmetric space
The graded algebra $\mathfrak{g}=\overbrace{\mathfrak{g}_{1}\oplus\mathfrak{g}_{0}}^{\mathfrak{h}}\oplus\overbrace{\mathfrak{g}_{-1}}^{\mathfrak{p}}\simeq\mathbb{R}^{*3,1}\oplus\overbrace{\mathfrak{co}(4,2)}^{\mathfrak{so}(3,1)\oplus\mathbb{R}}\oplus\mathbb{R}^{3,1}$ which is reductive, is also a symmetric space $\mathfrak{g}=\mathfrak{co}(4,2)\oplus\mathfrak{m}$ with $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}\simeq\mathbb{R}^{*3,1}\oplus\mathbb{R}^{3,1}$ with commutation relations, see (LABEL:eq:gradedLa) and [10]:
$[\mathfrak{g}_{0},\mathfrak{g}_{0}]\subset\mathfrak{g}_{0},\qquad[\mathfrak{g}_{0},\mathfrak{m}]\subset\mathfrak{m},\qquad[\mathfrak{m},\mathfrak{m}]\subset\mathfrak{g}_{0}\ .$ (1)
Let $\varphi=\varphi_{\mathfrak{g}_{1}}\oplus\varphi_{\mathfrak{g}_{0}}\oplus\varphi_{\mathfrak{g}_{-1}}=\bar{a}\oplus(c-k\mathds{1}_{4})\oplus b\in\mathfrak{g}$ with $a,b\in\mathbb{R}^{3,1}$, $c\in\mathfrak{so}(3,1)$, $k\in\mathbb{R}$ and $\bar{a}=\eta a^{T}$; $\bar{b}=\eta b^{T}\in\mathbb{R}^{*3,1}$. A matrix representation for $\mathfrak{g}$ is given by
$\varphi=\begin{pmatrix}k&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-k\end{pmatrix}=\begin{pmatrix}0&\bar{a}&0\\ 0&0&a\\ 0&0&0\end{pmatrix}+\begin{pmatrix}k&0&0\\ 0&c&0\\ 0&0&-k\end{pmatrix}+\begin{pmatrix}0&0&0\\ b&0&0\\ 0&\bar{b}&0\end{pmatrix}$ (14)
and the Lie bracket on $\mathfrak{g}$ is given by the commutator:
$[\varphi,\varphi^{\prime}]$ $=\begin{bmatrix}\begin{pmatrix}k&\bar{a}&0\\ b&c&a\\ 0&\bar{b}&-k\end{pmatrix},\begin{pmatrix}k^{\prime}&\bar{a}^{\prime}&0\\ b^{\prime}&c^{\prime}&a^{\prime}\\ 0&\bar{b}^{\prime}&-k^{\prime}\end{pmatrix}\end{bmatrix}$ (15)
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&k\bar{a}^{\prime}-k^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&\bar{a}a^{\prime}-\bar{a}^{\prime}a\\ k^{\prime}b-kb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-k^{\prime}a+ka^{\prime}\\ \bar{b}b^{\prime}-\bar{b}^{\prime}b&\bar{b}c^{\prime}-\bar{b}^{\prime}c-k\bar{b}^{\prime}+k^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
$=\begin{pmatrix}\bar{a}b^{\prime}-\bar{a}^{\prime}b&k\bar{a}^{\prime}-k^{\prime}\bar{a}+\bar{a}c^{\prime}-\bar{a}^{\prime}c&0\\ k^{\prime}b-kb^{\prime}+cb^{\prime}-c^{\prime}b&[c,c^{\prime}]+b\bar{a}^{\prime}-b^{\prime}\bar{a}+a\bar{b}^{\prime}-a^{\prime}\bar{b}&ca^{\prime}-c^{\prime}a-k^{\prime}a+ka^{\prime}\\ 0&\bar{b}c^{\prime}-\bar{b}^{\prime}c-k\bar{b}^{\prime}+k^{\prime}\bar{b}&\bar{b}a^{\prime}-\bar{b}^{\prime}a\end{pmatrix}$
A Cartan connection $\varpi$ with this $G/H$ Cartan geometry can be decomposed as :
$\varpi=\varpi_{\mathfrak{h}}\oplus\varpi_{\mathfrak{p}}=\overbrace{\varpi_{1}\oplus\varpi_{0}}^{\varpi_{\mathfrak{h}}}\oplus\varpi_{-1}=$ $\varpi_{1}^{a}\begin{pmatrix}0&\bar{e}_{a}&0\\ 0&0&e_{a}\\ 0&0&0\end{pmatrix}+\begin{pmatrix}\lambda&0&0\\ 0&A&0\\ 0&0&-\lambda\end{pmatrix}+\varpi_{-1}^{a}\begin{pmatrix}0&0&0\\ e_{a}&0&0\\ 0&\bar{e}_{a}&0\end{pmatrix}$ (25)
$=$ $\begin{pmatrix}\lambda&\dfrac{1}{\ell^{\prime}}\bar{\alpha}&0\\ \dfrac{1}{\ell}\beta&A&\dfrac{1}{\ell^{\prime}}\alpha\\ 0&\dfrac{1}{\ell}\bar{\beta}&-\lambda\end{pmatrix}$ (29)
where $A=\dfrac{1}{2}A^{ab}J_{ab}=\dfrac{1}{2}A^{ab}_{\mu}dx^{\mu}\otimes J_{ab}$ corresponds to the spin connection and $\beta=\beta^{a}e_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes e_{a}$ corresponds to the tetrad or soldering form used to define the metric of GR $\tilde{g}=\eta(\beta,\beta)$. $\alpha=\alpha^{a}e_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes e_{a}$ can be interpreted as a secondary tetrad. Finally $\lambda=\lambda_{\mu}dx^{\mu}$ is the $\mathbb{R}$-valued (dilations) part of the connection. Also : $\bar{\alpha}=\alpha^{a}\bar{e}_{a}=\alpha^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$ and $\bar{\beta}=\beta^{a}\bar{e}_{a}=\beta^{a}_{\mu}dx^{\mu}\otimes\bar{e}_{a}$. The expression of the curvature $\bar{\Omega}=d\varpi+\dfrac{1}{2}[\varpi,\varpi]$ of $\varpi$ is :
$\bar{\Omega}$ $=\bar{\Omega}_{\mathfrak{h}}\oplus\bar{\Omega}_{\mathfrak{p}}=\begin{pmatrix}f&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ \dfrac{1}{\ell}\Theta&F&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&\dfrac{1}{\ell}\bar{\Theta}&-f\end{pmatrix}=\begin{pmatrix}f&\dfrac{1}{\ell^{\prime}}\bar{\Pi}&0\\ 0&F&\dfrac{1}{\ell^{\prime}}\Pi\\ 0&0&-f\end{pmatrix}+\begin{pmatrix}0&0&0\\ \dfrac{1}{\ell}\Theta&0&0\\ 0&\dfrac{1}{\ell}\bar{\Theta}&0\end{pmatrix}$ (39)
$=\begin{pmatrix}d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta&\dfrac{1}{\ell^{\prime}}(d\bar{\alpha}+\lambda\wedge\bar{\alpha}+\bar{\alpha}\wedge A)&0\\ \dfrac{1}{\ell}(d\beta+\lambda\wedge\beta+A\wedge\beta)&dA+\dfrac{1}{2}[A,A]+\dfrac{1}{\ell\ell^{\prime}}(\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta})&\dfrac{1}{\ell^{\prime}}(d\alpha+A\wedge\alpha+\lambda\wedge\alpha)\\ 0&\dfrac{1}{\ell}(d\bar{\beta}+\bar{\beta}\wedge A+\lambda\wedge\bar{\beta})&-d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\beta}\wedge\alpha\end{pmatrix}$ (43)
with
$F=R+\dfrac{1}{\ell\ell^{\prime}}\phi\quad;\quad$ $\phi=\beta\wedge\bar{\alpha}+\alpha\wedge\bar{\beta}\quad;\quad f=d\lambda+\dfrac{1}{\ell\ell^{\prime}}\bar{\alpha}\wedge\beta$
$\Pi=d\alpha+A\wedge\alpha+\lambda\wedge\alpha\quad$ $\text{and}\quad\Theta=d\beta+\lambda\wedge\beta+A\wedge\beta=T+\lambda\wedge\beta\ .$
The Bianchi identities are given by $D\bar{\Omega}=d\bar{\Omega}+\varpi\wedge\bar{\Omega}-\bar{\Omega}\wedge\varpi=0$ which splits into the identities:
$(D\bar{\Omega})_{-1}=\dfrac{1}{\ell}(d\Theta-F\wedge\beta+\beta\wedge f+A\wedge\Theta-\Theta\wedge\lambda)=0$
$(D\bar{\Omega})_{\mathfrak{so}(3,1)}=\underbrace{dR+[A,R]}_{=0}+\dfrac{1}{\ell\ell^{\prime}}(d\phi+[A,\phi]+\beta\wedge\bar{\Pi}-\Pi\wedge\bar{\beta}+\alpha\wedge\bar{\Theta}-\Theta\wedge\bar{\alpha})=0$
$(D\bar{\Omega})_{\mathbb{R}}=-\mathds{1}\bigl{(}df+\dfrac{1}{\ell\ell^{\prime}}(\bar{\Theta}\wedge\alpha-\bar{\beta}\wedge\Pi)\bigr{)}=0$
$(D\bar{\Omega})_{1}=\dfrac{1}{\ell^{\prime}}(d\bar{\Pi}+\bar{\alpha}\wedge F-f\wedge\bar{\alpha}-\bar{\Pi}\wedge A+\lambda\wedge\bar{\Pi})=0.$
Remark.
If we impose a normal Cartan geometry such that $\bar{\Omega}_{\mathfrak{p}}=0$ and also consider a gauge where $\lambda=0$, then, as described in [7], $\bar{\Pi}$ and $\bar{\alpha}$ become respectively the Cotton and Schouten tensor while $F$ and $\Theta$ would correspond to the Weyl and torsion tensor. These tensors constitute the building blocks of conformal gravity.
3.2 Characteristic poylynomials for $G=SO(4,2)/\{\pm I\}$ and associated action
Recall the Levi-Civita symbols and tensors are still linked via the relations:
$\varepsilon_{abcd}=\varepsilon_{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;b}\beta^{-1\,\mu_{3}}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\mu_{4}}_{~{}~{}\,\,\;\;\;d}\quad\text{and}\quad\varepsilon^{abcd}=\varepsilon^{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{\,a}_{~{}\,\,\mu_{1}}\beta^{\,b}_{~{}\,\,\mu_{2}}\beta^{\,c}_{~{}\,\,\mu_{3}}\beta^{\,d}_{~{}\,\,\mu_{4}}.$
The trace $\operatorname{Tr}$ used to compute the Pontrjagin number associated to $\bar{\Omega}$ is given by
$\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi_{\mathfrak{so}}^{{}^{\prime}cd}+2\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}+2\eta_{ab}(\varphi_{-1}^{a}\varphi_{1}^{{}^{\prime}b}+\varphi_{1}^{a}\varphi_{-1}^{{}^{\prime}b}).$
Taking a linear combination of the Euler and Pontrjagin numbers ($\mathdutchcal{E}$ and $\mathdutchcal{P}$) associated respectively to $F$ and $\bar{\Omega}$ yields the action :
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})\Bigr{)}$ (1)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}F^{ab}\wedge F^{cd}-\dfrac{r}{8\pi^{2}}Tr(\bar{\Omega}\wedge\bar{\Omega})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\underbrace{\phi^{ab}\wedge\phi_{ab}}_{=0}-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2f\wedge f\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon_{abcd}\varepsilon^{r_{1}s_{1}cs_{2}}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r_{1}}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{4}}\beta^{\,\mu_{4}}_{~{}\,\,s_{2}}d^{4}x$
$\qquad\qquad\qquad\quad+\dfrac{16}{\ell^{2}\ell^{\prime 2}}\sqrt{|\tilde{g}|}\varepsilon_{abcd}\varepsilon^{as_{1}cs_{2}}\alpha^{~{}\,b}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{4}}\beta^{-1\,\mu_{4}}_{~{}~{}\,\,\;\;\;s_{2}}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{4}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{r_{1}s_{1}r_{2}s_{2}}R^{ab}_{\;\;\;r_{1}s_{1}}\beta^{\,c}_{~{}\,\,\mu_{1}}\alpha^{~{}\,d}_{~{}~{}\,\,\mu_{2}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;r_{2}}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;s_{2}}\eta_{ca}\eta_{db}d^{4}x$
$\qquad\qquad\quad\quad-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}.$
By considering the very particular case where [^4] This constraint does not seem to come from a mutation process as defined in Def.2.1. $\alpha=\dfrac{k}{2}\beta$, we get $\Pi=\dfrac{k}{2}\Theta=\dfrac{k}{2}(T+\lambda\wedge\beta)$, $\phi=\ell^{2}\xi$, $F=R+k\xi$ and one arrives at the following action:
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\mathscr{M})\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}(\dfrac{\bar{\Omega}_{\mathfrak{so}(3,1)}}{2\pi})+rP(\bar{\Omega})\Bigr{)}$ (2)
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\underbrace{\xi^{ab}\wedge\xi_{ab}}_{=0}-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}\Theta^{a}\wedge\Theta^{b}-2f^{2}\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi^{cd}+\dfrac{k^{2}\ell^{2}}{\ell^{\prime 2}}\xi^{ab}\wedge\xi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k\ell}{\ell^{\prime}}R^{ab}\wedge\xi_{ab}-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\bigl{(}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{8k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;b}d^{4}x+\dfrac{24k^{2}}{\ell^{2}\ell^{\prime 2}}\sqrt{|\tilde{g}|}d^{4}x\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell\ell^{\prime}}\sqrt{|\tilde{g}|}\varepsilon^{rs}_{~{}~{}ab}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;r}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;s}d^{4}x$
$\qquad\qquad\quad\quad-\dfrac{2k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge T^{b}-\dfrac{4k}{\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$
and the resulting action for $k=0$ is :
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$ (3)
which corresponds to the action (4) for $G=ISO(3,1)\Leftrightarrow k=0$, up to the $\lambda$ kinetic term. However, ignoring the boundary terms leaves the equations of motion unchanged if we keep the same matter Lagrangian as in the preceding example 2. While for $k\neq 0$ a direct comparison between $S^{\prime}$ (3) and $S_{G}$ (2) amounts to identifying the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{1}\\ \dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\alpha_{5}\\ \dfrac{ek^{2}}{8(\pi\ell\ell^{\prime})^{2}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{3}=-2\alpha_{5}\\ \alpha_{1}=-\dfrac{2ek}{r\ell\ell^{\prime}}\alpha_{3}\\ r=\dfrac{2\pi^{2}\ell\ell^{\prime}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (4)
The solution of (4) in terms of the parameters $e$, $r$, $k$, $\ell$ and $\ell^{\prime}$ is:
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell\ell^{\prime}}\alpha_{4}=\dfrac{2\ell\ell^{\prime}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{2k}{\ell\ell^{\prime}}\alpha_{3}=-\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r}\alpha_{2}\\ r=\dfrac{4\pi^{2}\ell\ell^{\prime}}{k}\alpha_{2}\end{cases}$ (5)
Hence, the resulting action reads:
$S_{G}[\varpi]=\int_{\mathscr{M}}$ $\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell\ell^{\prime}}(\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+rR^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi\ell\ell^{\prime})^{2}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad+\,\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}$
$\qquad\overbrace{-\,\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{Torsion-dilation~{}interaction}\overbrace{-\dfrac{r}{4\pi^{2}}d\lambda\wedge d\lambda}^{Dilation~{}kinetic~{}term}\Bigr{)}.$ (6)
Choosing the symmetric decomposition $\mathfrak{g}=\mathfrak{g}_{0}\oplus\mathfrak{m}$ with $\mathfrak{m}=\mathfrak{g}_{1}\oplus\mathfrak{g}_{-1}$ and $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$ such that the Killing metric is computed to be:
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})=\overbrace{-4\eta_{ac}\eta_{bd}\varphi_{\mathfrak{so}}^{ab}\varphi_{\mathfrak{so}}^{{}^{\prime}cd}+8\varphi_{\mathbb{R}}\varphi^{\prime}_{\mathbb{R}}}^{K_{0}}+\overbrace{8\eta_{ab}(\varphi_{-1}^{a}\varphi_{1}^{{}^{\prime}b}+\varphi_{1}^{a}\varphi_{-1}^{{}^{\prime}b})}^{K_{\mathfrak{m}}}$
and after fixing $\alpha=\dfrac{k}{2}\beta$, one gets for the metric:
$g(X,Y)=(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{8k}{\ell\ell^{\prime}}\,\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}.$
If one considers the usual covariant derivative $D_{\mathfrak{so}}=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$, the matter action similar to (2) is then given by :
$S_{M}[A,\beta]$ $=\dfrac{(\ell\ell^{\prime})^{2}}{384(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{so}}\psi\bigl{)}=ReS_{D}$ (7)
A more natural matter action $S^{\prime}_{M}$ with covariant derivative $D_{0}=d+\Phi_{\varepsilon}(\varpi_{0})=d+\Phi_{\varepsilon}(A-\lambda\mathds{1}_{4})=(\partial_{\mu}+\dfrac{1}{4}A_{ab,\mu}\gamma^{a}\gamma^{b}-3\lambda_{\mu})dx^{\mu}$ taking into account the full action of the Weyl algebra $\mathfrak{g}_{0}=\mathfrak{co}(3,2)$-valued part of the connection $\varpi_{0}$ is:
$S^{\prime}_{M}[A,\beta,\lambda]$ $=\dfrac{(\ell\ell^{\prime})^{2}}{384(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{0}\psi\bigl{)}$ (8)
$=\overbrace{S_{M}}^{ReS_{D}}-\dfrac{(\ell\ell^{\prime})^{2}}{128(2\ell^{\prime}+k\ell)k}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*(\lambda\psi)\bigl{)}.$ (9)
The study of the full action $S^{\prime}_{T}=S_{G}+S^{\prime}_{M}$ is left for a future work. In the following we will restrict ourselves to the study of the action $S_{T}=S_{G}+S_{M}$. For $k\neq 0$ the equations of motion for $\beta$, $A$ and $\lambda$ yield (neglecting boundary terms) :
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2R^{ab}\wedge\beta_{b}\delta_{ac}+T_{c}\wedge\lambda-\beta_{c}\wedge d\lambda)$
$=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta\beta^{c}}$ (10)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2T_{a}+\lambda\wedge\beta_{a})\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta A^{ab}}$ (11)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta,\lambda]}{\delta\lambda}$ $=-\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}T^{a}\wedge\beta_{a}=0=-\dfrac{\delta\mathscr{L}_{M}[A,\beta,\lambda]}{\delta\lambda}\ .$ (12)
Equation (12) implies either $r=0$ or $T=0$ since the definition of a Cartan connection requires $\beta$ to be an isomorphism. For $r=0$, one then recovers the usual equations of motion associated to Einstein-Cartan gravity, with the torsion $T$ linked to the spin density of matter. In the other case $T=0$, the equations of motion become:
$\dfrac{ek}{4\pi^{2}\ell\ell^{\prime}}(R^{ab}+\dfrac{2k}{\ell\ell^{\prime}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}(2R^{ab}\wedge\beta_{b}\delta_{ac}-\beta_{c}\wedge d\lambda)$ $=\tau_{c}$ (13)
$\dfrac{rk}{2\pi^{2}\ell\ell^{\prime}}\lambda\wedge\beta_{a}\wedge\beta_{b}$ $=\dfrac{1}{2}\mathfrak{s}_{ab}$ (14)
$T$ $=0$ (15)
where equation (14) links spin density to the scalar field:
$\lambda_{c}=\lambda_{\mu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}=\dfrac{\pi^{2}\ell\ell^{\prime}}{12rk}\mathfrak{s}^{ab}_{~{}~{}abc}\ .$ (16)
Using the relation $\gamma=\dfrac{2\alpha_{1}}{2\alpha_{2}-\alpha_{5}}$, given in [14, p13], we can identify $\alpha_{5}=-\alpha_{2}=-\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=-\dfrac{2\alpha_{1}}{3\gamma}$, with $\gamma$ the Barbero-Immirzi parameter. Provided we set $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$, and applying the Hodge star *** to the first equation and replacing $\lambda$ according to (16) gives us the Einstein equations modified by the Barbero-Immirzi part of the action with bare cosmological constant $\Lambda_{0}$ and with matter sources given by the energy-momentum tensor and a particular combination of derivatives of the spin density of matter (the latter resulting from the presence of the dilation 1-form $\lambda$ of the connection):
$G_{kc}+\dfrac{2}{3\gamma}R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda_{0}\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{ck}-\dfrac{1}{48}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}).$ (17)
Separating the energy-momentum tensor $\tau_{ck}$ as $\tau_{ck}=\tau_{M,ck}-\rho_{\text{vac}}\eta_{ck}$ with $-\rho_{\text{vac}}\eta_{ck}$ the part related to the vacuum energy density $\rho_{\text{vac}}\geq 0$ then leads to:
$G_{kc}+\dfrac{2}{3\gamma}R^{ab}_{~{}~{}rs}\varepsilon^{rs}_{~{}~{}bk}\delta_{ac}+\Lambda\eta_{kc}=-\dfrac{1}{2\alpha_{1}}(\tau_{M,ck}-\dfrac{1}{48}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck})$ (18)
with $\Lambda=\Lambda_{0}+\Lambda_{\text{vac}}=\overbrace{-\dfrac{6k}{\ell\ell^{\prime}}}^{\Lambda_{0}}\overbrace{-\dfrac{\rho_{\text{vac}}}{2\alpha_{1}}}^{\Lambda_{\text{vac}}}=-\dfrac{6k}{\ell\ell^{\prime}}-\dfrac{2\pi^{2}\ell\ell^{\prime}\rho_{\text{vac}}}{ek}=\Lambda_{0}+8\pi G\rho_{\text{vac}}$ the effective cosmological constant, and $\Lambda_{V}$ the vacuum contribution to the cosmological constant. One can infer that changes in cosmological history of the rate of expansion of the universe could be partly linked to spin-density variations in this way.
Remark.
In this particular case the fact that Torsion is null implies that the Einstein tensor $G_{kc}$ is symmetric (see [15, p3] for more details). When taking into account that $\eta_{kc}$ is symmetric as well, the anti-symmetricity of $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}$ implies that the energy-momentum tensor $\tau_{ck}$ sports an anti-symmetric part if $\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;ck}\neq 0$. This confirms that we can consider a matter Lagrangian that depends on the spin connection $A$ as we explicitly did.
Replacing $e=\dfrac{8\pi^{2}\alpha_{1}\ell\ell^{\prime}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell\ell^{\prime}=-\dfrac{6k}{\Lambda_{0}}$ in the action and $\dfrac{rk}{4\pi^{2}\ell\ell^{\prime}}=\dfrac{2\alpha_{1}}{3\gamma}$, we obtain:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}$
$\qquad\qquad\overbrace{-\dfrac{4\alpha_{1}}{3\gamma}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}+\overbrace{\underbrace{\dfrac{4\alpha_{1}}{\Lambda_{0}\gamma}}_{-2\vartheta}d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}\biggr{)}.$ (19)
Remark.
The interpretation with respect to Chern-Simons modified gravity remains, only the potential term changes:
$V(\vartheta)$ $=\dfrac{2\alpha_{1}}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}$
$\qquad\qquad\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\overbrace{-\dfrac{4\alpha_{1}}{3\gamma}\eta_{ab}T^{a}\wedge\lambda\wedge\beta^{b}}^{\text{ Dilation-Torsion interaction}}-\overbrace{2\vartheta d\lambda\wedge d\lambda}^{\text{ Dilation kinetic term }}$ (20)
Setting $\lambda=0$ we retrieve the action (15), namely:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\biggl{(}\overbrace{\alpha_{1}(R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{2}{3\gamma}R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Holst}+\overbrace{\dfrac{\Lambda_{0}\alpha_{1}}{6}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\qquad+\overbrace{\dfrac{3\alpha_{1}}{2\Lambda_{0}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}+\overbrace{\underbrace{\dfrac{2\alpha_{1}}{\Lambda_{0}\gamma}}_{\vartheta}R^{ab}\wedge R_{ab}}^{Pontrjagin}\overbrace{-\dfrac{2\alpha_{1}}{3\gamma}\bigl{(}T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b}\bigr{)}}^{Nieh-Yan}\biggr{)}\Biggr{)}.$
This shows that by imposing specific constraints on $\alpha$ and $\lambda$ for $G=SO(4,2)/\{\pm I\}$, one can recover the characteristic polynomial based action for $G=SO(4,1),~{}SO(3,2),~{}\text{or }ISO(3,1)$ with respective constraints:
$\left\{\begin{array}[]{ll}&\lambda=0\text{ and }\alpha=\dfrac{k}{2}\beta\\ &\left\{\begin{array}[]{ll}sign(k\ell\ell^{\prime})=-,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}>0&\Leftrightarrow G=SO(4,1)\\ k=0&\Leftrightarrow G=ISO(3,1)\\ sign(k\ell\ell^{\prime})=+,\quad\Lambda_{0}=-\dfrac{6k}{\ell\ell^{\prime}}<0&\Leftrightarrow G=SO(3,2)\end{array}\right.\end{array}\right.$ (26)
Remark.
$\alpha_{2}=0$ and $\lambda=0$ give different actions but the same equations of motion if we neglect the boundary terms.
In the case $\alpha\neq\dfrac{k}{2}\beta$ ($\alpha$ is related to the special conformal transformations) it is interesting to notice that the Palatini action and the cosmological constant term in action (15) respectively become the terms proportional to $\varepsilon_{abcd}R^{ab}\wedge\phi^{cd}$ and $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ (it is recalled that $\phi$ was defined in eq.(44)) in the action:
$S[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bigl{(}R^{ab}\wedge R^{cd}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi^{cd}+\dfrac{1}{(\ell\ell^{\prime})^{2}}\phi^{ab}\wedge\phi^{cd}\bigr{)}$
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2}{\ell\ell^{\prime}}R^{ab}\wedge\phi_{ab}-\dfrac{4}{\ell\ell^{\prime}}\eta_{ab}\Pi^{a}\wedge\Theta^{b}-2d\lambda\wedge d\lambda\bigr{)}\Bigr{)}$ (27)
where the fact that $\alpha^{~{}\,a}_{~{}~{}\,\,\mu}\beta^{\,\mu}_{~{}\,\,b}\neq\dfrac{k}{2}\delta^{a}_{b}$ turns out to be an obstruction for $\varepsilon_{abcd}\phi^{ab}\wedge\phi^{cd}$ to be a constant volume form.
4 Conclusion
We have shown that a large class of models for gravitation, including the standard theory of general relativity, can be solely derived from topological considerations. This has been achieved by considering an action for gravitation as a linear combination of the Euler and Pontrjagin numbers of a manifold $\mathscr{M}$ with specific underlying $G/H$ Cartan geometry. Since the action is expressed as a linear combination of characteristic numbers constructed with various parts of the curvature of a Cartan connection, the action is invariant under the choice of such a Cartan connection up to the integral over $\mathscr{M}$ of an exact term. Two classes of examples have been studied from which Einstein’s equations (with positive or negative bare cosmological constant $\Lambda_{0}$) can be obtained upon making a restrictive choice (no Pontrjagin number). The first example corresponds to the case of a $\mathfrak{g}=\left\{\begin{array}[]{ll}\mathfrak{so}(4,1)\text{ for }\Lambda_{0}>0\\ \mathfrak{so}(3,2)\text{ for }\Lambda_{0}<0\end{array}\right.$-valued Cartan connection $\varpi$ defined on the $H=SO(3,1)$-principal-bundle $\mathscr{P}$. The computed action gives the Holst action together with the Euler, Pontrjagin and Nieh-Yan topological terms, see (15). The coupling constants of these topological terms are entirely determined by the two parameters of the Holst action ($G$ and $\gamma$) and the bare cosmological constant $\Lambda_{0}$. The second example deals with the Möbius geometry. In this case a very peculiar choice of the secondary tetrad $\alpha$ within the $\mathfrak{g}=\mathfrak{so}(4,2)$-valued Cartan connection $\varpi$ defined on the $H=CO(4,2)\ltimes\mathbb{R}^{3,1}$-principal-bundle $\mathscr{P}$ yields, either the same results as in the previous class of examples if one drops out dilations, or the Einstein’s equations with an additional source term for the curvature depending on a particular variation of spin density $-\dfrac{\pi G}{6}\partial_{\mu}\mathfrak{s}^{ab}_{~{}~{}ab\nu}\varepsilon^{\mu\nu}_{~{}~{}\;kc}$ when coupled to matter. In these two classes of examples treated in the paper, the action as linear combination of the Euler and Pontrjagin numbers contains the Holst action showing that the Holst action may have a topological origin. This opens the possibility to figure out a topological formulation of LQG.